The width of a rectangular tank is X cm. It's height is 2/3 it's width and it's length is four times it's height. The volume of the tank in litres is


1
Expert's answer
2021-04-29T12:06:23-0400

Solution.

A rectangular tank is a generalized form of a cube, where the sides can have varied lengths. It is bounded by six faces, three of which meet at its vertices, and all of which are perpendicular to their respective adjacent faces. The equation for calculating the volume of a rectangle is:

volume= length × width × height.

Let be "x" cm is the width, then "\frac{2}{3}x" cm is the height and "\frac{8}{3}x" cm is the length.

Find volume:


"x\cdot \frac{2}{3}x \cdot \frac{8}{3}x =\frac{16}{9}x^3 (cm^3)."

Answer. "\frac{16}{9}x^3."